OpenStudy (anonymous):

@waterineyes A painter uses 2/3 gallon of paint on one job and 3/5 of the same paint on another. How many gallons of paint did the painter use for both jobs? A. 1/15 B. 1 1/3 C. 1 1/15 D. 1 4/15

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

Add both..

OpenStudy (anonymous):

its was D

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

\[\frac{19}{15} = 1 \frac{4}{19}\]

OpenStudy (anonymous):

sorry 15..

OpenStudy (anonymous):

\[\frac{19}{15} = 1 \frac{4}{15}\]